OpenStudy (anonymous):

MENTAL ABILITY find the missing term 4,13,53,160,641,? a-1923 b-1282 c-2564 d-none of these and tell me how you identify the pattern pls!!! help

OpenStudy (zarkon):

d

OpenStudy (anonymous):

ok well here is the pattern i saw... 13 = (4x3) + 1 53 = (13x4) + 1 160 = (53 x 3) + 1 641 = (160 x 4) + 1 so the next term should be (641x3) + 1 = 1923 + 1 = 1924 but this is not here so I must be wrong....

OpenStudy (anonymous):

anyone care to explain?

OpenStudy (akshay_budhkar):

u r right.. so ur answer is d

OpenStudy (anonymous):

oh yer sorry i'm not too good at seeing words lol

OpenStudy (anonymous):

there is another question which i am unable to understand

OpenStudy (anonymous):

38th question

OpenStudy (anonymous):

and explain how do you find the pattern

OpenStudy (anonymous):

oops it alternates 3x+1 4x+1 3x+1 4x+1

OpenStudy (anonymous):

it's 22

OpenStudy (anonymous):

how

OpenStudy (anonymous):

You have to look at each row o the grid. The difference between 39 and 37 is 2. The difference between 28 and 23 is 5, the difference between 38 and 42 is 4. It appears as though these differences are being multipled by 11...

OpenStudy (anonymous):

but one of them is -4 the others are +5, +2 ?

OpenStudy (anonymous):

yeah I see what you mean, but I didn't assume an implicit logic to how the differences are calculated...I will keep looking at it! :D It may not matter. Could be something like "the largest number in row 1 minus the smallest number in row 1 multiplied by 11 gives the middle number of row 1".

OpenStudy (akshay_budhkar):

but is 22 there in the 4 option below, then this explanation may be accepted.. i m still unsure if this is the right answer, same doubt as Broken Fixer

OpenStudy (anonymous):

22 is option (A)

OpenStudy (akshay_budhkar):

oh ok i didnt scroll down my bad

OpenStudy (anonymous):

thanks every body

OpenStudy (anonymous):

also the answer to question 37 is C not A (the one you ticked). The first letter in each string of three corresponds to the sequence {3, 6, 9, 12, 15}. The second letter in each string of three corresponds to the sequence {3-1, 6-1, 12-1, 15-1}. The third letter in each string of three corresponds to the sequence {3-2, 6-2, 9-2, 12-2, 15-2}. So the next item in your sequence is ONM.